OpenStudy (anonymous):

A 50 N block rests on an adjustable inclined plane; the maximum angle just before it slips is 28 degrees with the horizontal. What is the coefficient of static friction between the block and the plane surfaces?

OpenStudy (anonymous):

Well, call the angle alpha: The point when it slips is when the maximum frictional force and the component for gravity pulling the block down the slope are equal. \[(1)mg \sin(\alpha) - F_{Fmax}=0\] From the definition of static friction coefficient: \[(2)\mu_{static} \times N = F_{Fmax}\] N, the normal force, can be rewritten: \[(3)N= mg \cos(\alpha)\] substituting (3) into (2): \[(4) \mu_{static} \times mg \cos (\alpha) = F_{Fmax}\] And (4) into (1): \[(5)mg \sin(\alpha)- \mu_{static}mgcos(\alpha)=0\] Divide through by mg: \[(6)\sin(\alpha)-\mu_{static} \cos(\alpha)=0\] Rearrange: \[(7)\mu_{static}=\tan(\alpha)\] So... \[\alpha \approx 0.531\] An interesting result is that the friction coefficient is independent of the mass or surface area of the object, the 50N was unnecessary.

OpenStudy (anonymous):

Sorry, that is supposed to be the friction coefficient above, not alpha. Mixing up my Greek alphabet